# image generation; decoupled cross-attention; latent diffusion model; controllability

### v3

@@ Abstract
The creation of landscape paintings is a meticulous and skill-intensive endeavor, demanding exceptional artistic proficiency. With the development of artificial intelligence (AI) diffusion models, using these models for generating landscape paintings has become a novel method for addressing the aforementioned challenges. Herein, we build a unique traditional landscape painting dataset. Besides, we propose a high-fidelity, controllable model for landscape painting generation (LPGen), pioneering a novel multi-modal framework by incorporating the structure controller and the style controller into the diffusion model. Specifically, the structure controller module utilizes the Decoupled Cross-Attention technique to separately manage the structural information of an image across different domains, allowing for precise control and regulation of specific elements and attributes in the generated image. Meanwhile, the style controller dynamically adjusts the features of the input image, enabling the generative model to accurately capture and reflect the style of the canny image. The research results indicate that the proposed method outperforms existing approaches in landscape painting generation and exceeds the current state-of-the-art. Additionally, the proposed LPGen network excels in controlling the structure and style of landscape paintings, producing highly accurate images, and fostering further research in deep learning-based landscape painting generation.
keywords:
image generation; decoupled cross-attention; latent diffusion model; controllability
@@
[…]
1 Introduction
Landscape painting holds a significant place in traditional art, depicting natural scenes to convey the artist’s aesthetic vision and emotional connection to nature Yao et al. (2024); Maggos et al. (2019); Chakhim et al. (2020); Asperti et al. (2023); Wang et al. (2023); Zengguo Sun (2024). This genre utilizes tools such as ink, water, paper, and brushes, emphasizing artistic conception, brush and ink intensity variations, balanced composition, and effective use of negative space Zengguo Sun (2024); Teresa Kurkiewicz (2023). With the development of computer technology, especially diffusion, the generation of landscape paintings through modern algorithms has become a promising area of research Yao et al. (2024); Lin et al. (2018); Sun et al. (2024).
The creative process of landscape painting involves several intricate stages: outlining, chapping, rubbing, moss-dotting, and coloring, as shown in Figure 1 (a) Zengguo Sun (2024). Outlining is the initial stage where the artist sketches the basic structure and composition of the landscape Zengguo Sun (2024); Yao et al. (2024); Teresa Kurkiewicz (2023). Chapping involves the careful depiction of textures, requiring an understanding of how to use brushstrokes to convey different natural surfaces Teresa Kurkiewicz (2023). Rubbing and dyeing necessitate meticulous control over ink intensity and moisture, balancing between depth and subtlety in shading Zengguo Sun (2024). Moss-dotting requires a flexible and natural technique to create the illusion of texture and dimension, adding complexity and life to the painting Teresa Kurkiewicz (2023). Finally, coloring demands a keen sense of color harmony to ensure the overall unity and aesthetic appeal of the artwork Yao et al. (2024). The complexity and precision involved in each stage contribute to the overall challenge of creating landscape paintings Zengguo Sun (2024); Yao et al. (2024); Teresa Kurkiewicz (2023).
Recent technological advancements, particularly in deep learning, have empowered computers to emulate landscape paintings’ artistic styles and brushwork. These technologies can generate artworks that capture the rich cultural essence of China, preserving traditional art and opening new paths for creative expression Yang and Hu (2022); Li (2024); Way et al. (2023); Sun et al. (2023). This technological innovation supports the preservation and evolution of conventional culture Way et al. (2023); Sun et al. (2023); Way et al. (2023). Research into the generation of landscape paintings is primarily divided into traditional and deep learning-based methods. Traditional methods include non-photorealistic rendering, image-based approaches, and computer-generated simulations. Non-photorealistic rendering uses computer graphics to simulate various artistic effects, such as ink diffusion, through specialized algorithms Xin Zhang (2022); Andreas Semmo (2022). Image-based methods involve collecting brush stroke texture primitives (BSTP) from hand-drawn samples and employing them to map multiple layers to create mountain imagery Junho Kim (2024); Lu and Finkelstein (2023). Despite their sophistication, the quality of computer-generated simulations often falls short Xin Zhang (2022); Grubert et al. (2023).
@@ 1 Introduction
Currently, many researchers are using generative adversarial networks (GANs) to generate landscape paintings Park et al. (2021); Tang et al. (2022); Lee and Kang (2022); Zhang et al. (2023). However, due to the distinctive artistic style of landscape painting, GAN-based methods often face challenges such as unclear outlines, uncontrollable details, and inconsistent style migration Yang and Huang (2022). By visual generation, such as stable diffusion (SD), the artistic effects of landscape painting can be effectively simulated and reproduced Rombach et al. (2022); Liang and Zhao (2023). Diffusion models (DMs) have been successfully applied across various domains, including clothing swapping, face generation, dancing, and anime creation, demonstrating their versatility in generating high-quality realistic content from diverse inputs Ho et al. (2020); Meng and Zhang (2023); Liu and Chen (2023). For instance, diffusion models can produce highly realistic images nearly indistinguishable from actual photographs in face generation Choi and Lee (2023); Yang and Liu (2023). Similarly, in anime creation, they can generate detailed and stylistically consistent images from simple prompts Liu and Chen (2023); Park and Kim (2023). Although this method preserves the stylistic features of traditional landscape paintings, it does not allow precise control over the structure and style Liang and Zhao (2023); Rombach et al. (2022).
@@
2 Related Work
2.1 GAN-Based Method
@@
Due to the rapid advancements in deep learning, significant progress has been made in various visual tasks Shen et al. (2023a, b). Based on the input content, landscape painting generation methods can be categorized into three types: nothing-to-image, image-to-image, and text-to-image. For the first type, it does not require any input information to generate landscape paintings. An example is the Sketch-And-Paint GAN (SAPGAN) Xue (2021), the first neural network model capable of automatically generating traditional landscape paintings from start to finish. SAPGAN consists of two GANs: SketchGAN for generating edge maps and PaintGAN for converting these edge maps into paintings. Another example is an automated creation system Luo et al. (2022) based on GANs for landscape paintings, which comprises three cascading modules: generation, scaling, and super-resolution.
@@
For the second type, a photo is input for style transfer. An interactive generation method Zhou et al. (2019) allows users to create landscape paintings by simply outlining with essential lines, which are then processed through a recurring adversarial generative network (RGAN) model to generate the final image. Another approach, neural abstraction style transfer Li et al. (2019), leverages the MXDoG filter and three fully differentiable loss terms. The ChipGAN architecture Sun et al. (2022) is an end-to-end GAN model that addresses critical techniques in ink painting, such as blanks, brush strokes, ink tone and spread. Furthermore, an attentional wavelet network Sun et al. (2024) utilizes wavelets to capture high-level mood and local details of paintings via the 2-D Haar wavelet transform. Lastly, the Paint-CUT model Wang et al. (2023) intelligently creates landscape paintings by utilizing a shuffle attentional residual block and edge enhancement techniques.
@@
For the third type, the output image is guided by textual input information. A novel system Yuan et al. (2023) for Polaca transforms classical poetry into corresponding artistic landscape paintings and calligraphic works. Another approach, controlled landscape painting generation (CCLAP) Wang et al. (2023), is based on the latent diffusion model (LDM) Rombach et al. (2022) and comprises two sequential modules: a content generator and a style aggregator. Although this method allows text-based control of the image, it still faces challenges related to the randomness of image generation and limited controllability.
2.2 Diffusion-Based Method
@@
DM Ho et al. (2020) is a deep learning framework primarily utilized for image processing and computer vision tasks. Fundamentally, DM work by corrupting the training data by continuously adding Gaussian noise, and then learning to recover the data by reversing this noise process. DM can generate detailed images from textual descriptions and are also effective for image restoration, image drawing, text-to-image, and image-to-image tasks Lin et al. (2023); Zhang et al. (2024a, b); Fei et al. (2023). Essentially, by providing a textual description of the desired image, SD can produce a realistic image that aligns with the given description Ho et al. (2020). These models reframe the "image generation" process into a "diffusion" process that incrementally removes noise. Starting with random Gaussian noise, the models progressively eliminate it through training until it is entirely removed, yielding an image that closely matches the text description Ho et al. (2020); Zhang et al. (2024a). However, a significant drawback of this approach is the considerable time and memory required, especially for high-resolution image generation Lin et al. (2023). LDM Rombach et al. (2022) was developed to address these limitations by significantly reducing memory and computational costs. This is achieved by applying the diffusion process within a lower-dimensional latent space instead of the high-dimensional pixel space Lin et al. (2023).
@@
The emergence of large text-to-image models that can create visually appealing images from brief descriptive prompts has highlighted the remarkable potential of AI. However, these models encounter challenges such as uneven data availability across specific domains compared to the generalized image-to-text domain. In addition, some tasks require more precise control and guidance than simple prompts can provide. ControlNet Zhang et al. (2023) addresses these issues by generating high-quality images based on user-provided cues and controls, which can fine-tune performance for specific tasks. Meanwhile, IP-Adapter Ye et al. (2023) employs decoupled cross-attention mechanisms for the characteristics of the text and the image. Inspired by these approaches, our study incorporates ControlNet’s additional structural controls and the IP-Adapter’s style guiding the generation of landscape paintings.
3 Proposed Method
@@ 3.4 Training and Inference
$L_{\text{simple}}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t}\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\big{(}\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t\big{)}\|^{2}.$ (3)
@@
$\begin{split}\hat{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)=w\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{i},t)+(1-w)\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t).\end{split}$ (4)
@@ 3.5 Datasets
Data Collection. Datasets serve as the foundation for the rapid advancement of artificial intelligence, with a few open datasets significantly contributing to this progress. For example, the traditional landscape painting dataset  Xue (2021) is the only accessible dataset, but it has limitations in terms of data volume and lacks categorization or identification of the image data. To address this problem, we firstly have curated a collection of digital paintings sourced through websites such as Baidu, artwork websites, photographs from landscape painting books, and digital museum databases. In the following, we enlisted the help of professors and landscape artists to assist in the classification process. All the landscape paintings are classified into three main categories: azure green landscape, ink wash landscape, and light vermilion landscape.
@@
Data Preprocessing. Since the training data for this study should be in $512\times 512$ format, preprocessing the cleaned data is an essential step. First, each image was scaled so that its shorter side was 512 pixels, maintaining the original aspect ratio. For images with an aspect ratio less than 1.5, a $512\times 512$ section was cropped from the center and saved. For images with an aspect ratio greater than 1.5, after cropping the $512\times 512$ section from the center, the center point was moved 256 pixels in both directions along the longer side to crop and save two additional images. If the cropped image was not square, the process was halted, and the current cropped image was discarded. The distribution of the number of pictures is shown in Table 1.
Table 1: Landscape painting type distribution of three distinct styles in our collected dataset. The data comes from Harvard, Smithsonian, Metropolitan, Baidu, and Princeton.
@@
Data Pairing. The data format for this experiment is image-text pairs because it allows the model to learn the association between visual content and descriptive language, enhancing its ability to generate images that align with specific textual descriptions, thereby improving the overall accuracy and relevance of the generated images. Although image-text pairs previously required manual annotation, we now utilize BLIP-2 to automatically generate the corresponding text information for the collected images. Although the text generated by BLIP-2  Li et al. (2022) may be inaccurate, we invited artists who curated descriptive texts specific to landscape paintings. To increase the diversity of text descriptions, we provided multiple expressions for each text and specified the painting type within the descriptions. Finally, we saved the image paths and corresponding text descriptions in JSON files.
3.6 Evaluation Metrics
Several metrics are commonly used to assess how closely generated images resemble reference images in evaluating image and structural similarity. This document explains the meanings, formulas and applications of six key metrics: the learned perceptual image patch similarity (LPIPS) Zhang et al. (2018), gram matrix Gatys et al. (2015), histogram similarity Swain and Ballard (1991), chamfer match score Barrow and Tenenbaum (1977), hausdorff distance Huttenlocher et al. (1993), and contour match score Belongie et al. (2002).
@@ 3.6 Evaluation Metrics
Histogram Similarity (Bhattacharyya Distance). The histogram similarity is a metric for evaluating the similarity of color distributions of two images and is widely used in image processing and computer vision. By comparing the color histograms of the images, the similarity of the images in terms of color can be determined.The bhattacharyya Distance  Bhattacharyya (1943) is a commonly used histogram similarity metric, especially for similarity calculation of probability distributions.
@@
4 Experiment and Analysis
4.1 Generative Controllability
@@
4.2 Qualitative Analysis
To accurately and objectively assess the effectiveness of the proposed LPGen model in generating high-quality landscape paintings, we conduct a comprehensive comparative analysis against several current state-of-the-art methods, including Reference Only Zhang et al. (2023), Double ControlNet  Zhang et al. (2023), and Lora  Hu et al. (2021).
@@ 4.3 Quantitative analysis
Table 2: Quantitative comparison of the proposed LPGen with several state-of-the-art models. LPIPS evaluates perceptual similarity, gram matrix (GM) measures texture correlations for style similarity, and histogram similarity (HS) assesses color distribution. chamfer match score (CMS i) focuses on average edge similarity, hausdorff distance (HD) emphasizes worst-case edge similarity, and contour match score (CMS ii) evaluates overall shape similarity.
{adjustwidth}
[…]
@@
The landscape paintings generated by the proposed method were quantitatively compared with those generated by Reference Only, Double ControlNet, and Lora. Each model generated 15 images for each pair of style and structure, resulting in a total of 60 landscape painting images. For each model, the highest values for six metrics of the generated images were recorded: gram matrix similarity, histogram similarity, LPIPS, chamfer match score, hausdorff distance, and contour match score. The scores of the different diffusion models are shown in Table 2.
@@
In Table 2, the first three metrics, the gram matrix, histogram similarity, and LPIPS, are used to analyze the structural similarity between model outputs and reference images. The table shows that LPGen performs best with respect to the gram matrix, suggesting that LPGen’s generated images have the most similar texture style to the reference image. Furthermore, LPGen performs best with respect to histogram similarity, suggesting that LPGen-generated images have the most similar color distribution to the reference image. For structural similarity, LPGen excels in both gram matrix and histogram similarity, while Lora performs best in LPIPS due to the structure controller. To repetitive fine-tuning, Lora performs best in terms of LPIPS, indicating that Lora’s generated images are most perceptually similar to the reference image. Taking these factors into account, LPGen is the best model for overall structural Tsimilarity.
@@
In Table 2, the last three metrics, the chamfer match score, hausdorff distance, and contour match score, are used to analyze the style similarity between different model outputs and the reference image, mainly focusing on edges and contours. Obviously, LPGen performs best with respect to the chamfer match score, suggesting that LPGen’s generated images have the most similar edges to the reference image. Meanwhile, LPGen performs best with respect to the hausdorff distance, suggesting that LPGen’s generated images are closest to the reference image regarding edge similarity. Furthermore, LPGen performs best with respect to the contour match score, indicating that LPGen’s generated images have the most similar contour shapes to the reference image. Considered comprehensively, the landscape paintings generated by the proposed model have the most similar edges and contour shapes to the reference image, making it the best model for style similarity.
4.4 Visual Assessment

### v4

@@ Abstract
Generating high-fidelity landscape paintings remains a challenging task that requires precise control over both structure and style. In this paper, we present LPGen, a novel diffusion-based model specifically designed for landscape painting generation. LPGen introduces a decoupled cross-attention mechanism that independently processes structural and stylistic features, effectively mimicking the layered approach of traditional painting techniques. Additionally, LPGen proposes a structural controller, a multi-scale encoder designed to control the layout of landscape paintings, striking a balance between aesthetics and composition. Besides, the model is pre-trained on a curated dataset of high-resolution landscape images, categorized by distinct artistic styles, and then fine-tuned to ensure detailed and consistent output. Through extensive evaluations, LPGen demonstrates superior performance in producing paintings that are not only structurally accurate but also stylistically coherent, surpassing current state-of-the-art models. This work advances AI-generated art and offers new avenues for exploring the intersection of technology and traditional artistic practices. Our code, dataset, and model weights will be publicly available.
Index Terms:
image generation, decoupled cross-attention, latent diffusion model, controllability.
@@
=-21pt
I Introduction
Landscape painting holds a significant place in traditional art, depicting natural scenes to convey the artist’s aesthetic vision and emotional connection to nature [1, 2, 3, 4, 5]. This genre utilizes tools such as ink, water, paper, and brushes, emphasizing artistic conception, brush and ink intensity variations, balanced composition, and effective use of negative space [8, 6]. With the development of computer technology, especially diffusion, the generation of landscape paintings through modern algorithms has become a promising area of research [6].
The creative process of landscape painting involves several intricate stages: outlining, chapping, rubbing, moss-dotting, and coloring, as shown in Figure 1 (a) [6]. Outlining is the initial stage where the artist sketches the basic structure and composition of the landscape [6, 7]. Chapping involves the careful depiction of textures, requiring an understanding of how to use brushstrokes to convey different natural surfaces. Rubbing and dyeing necessitate meticulous control over ink intensity and moisture, balancing between depth and subtlety in shading [6]. Moss-dotting requires a flexible and natural technique to create the illusion of texture and dimension, adding complexity and life to the painting. Finally, coloring demands a keen sense of color harmony to ensure the overall unity and aesthetic appeal of the artwork [7]. The complexity and precision involved in each stage contribute to the overall challenge of creating landscape paintings [6, 7].
Recent technological advancements, particularly in deep learning, have empowered computers to emulate landscape paintings’ artistic styles and brushwork. These technologies can generate artworks that capture the rich cultural essence of China, preserving traditional art and opening new paths for creative expression [8, 9, 10, 11]. This technological innovation supports the preservation and evolution of conventional culture [10, 11, 12]. Research into the generation of landscape paintings is primarily divided into traditional and deep learning-based methods. Traditional methods include non-photorealistic rendering, image-based approaches, and computer-generated simulations. Non-photorealistic rendering uses computer graphics to simulate various artistic effects, such as ink diffusion, through specialized algorithms [13, 14]. Image-based methods involve collecting brush stroke texture primitives (BSTP) from hand-drawn samples and employing them to map multiple layers to create mountain imagery [15, 16]. Despite their sophistication, the quality of computer-generated simulations often falls short [13].
@@ I Introduction
Currently, many researchers are using generative adversarial networks (GANs) to generate landscape paintings [17, 18, 19]. However, due to the distinctive artistic style of landscape painting, GAN-based methods often face challenges such as unclear outlines, uncontrollable details, and inconsistent style migration [20]. By visual generation, such as stable diffusion (SD), the artistic effects of landscape painting can be effectively simulated and reproduced [21, 22]. Diffusion models (DMs) have been successfully applied across various domains, including clothing swapping, face generation, dancing, and anime creation, demonstrating their versatility in generating high-quality realistic content from diverse inputs [23, 24, 25]. For instance, diffusion models can produce highly realistic images nearly indistinguishable from actual photographs in face generation [26, 27]. Similarly, in anime creation, they can generate detailed and stylistically consistent images from simple prompts [28]. Although this method preserves the stylistic features of traditional landscape paintings, it does not allow precise control over the structure and style [22, 21].
@@
II Related Work
II-A GAN-Based Method
@@
Due to the rapid advancements in deep learning, significant progress has been made in various visual tasks [29, 30]. Based on the input content, landscape painting generation methods can be categorized into three types: nothing-to-image, image-to-image, and text-to-image. For the first type, it does not require any input information to generate landscape paintings. An example is the Sketch-And-Paint GAN (SAPGAN) [31], the first neural network model capable of automatically generating traditional landscape paintings from start to finish. SAPGAN consists of two GANs: SketchGAN for generating edge maps and PaintGAN for converting these edge maps into paintings. Another example is an automated creation system [32] based on GANs for landscape paintings, which comprises three cascading modules: generation, scaling, and super-resolution.
@@
For the second type, a photo is input for style transfer. An interactive generation method [33] allows users to create landscape paintings by simply outlining with essential lines, which are then processed through a recurring adversarial generative network (RGAN) model to generate the final image. Another approach, neural abstraction style transfer [34], leverages the MXDoG filter and three fully differentiable loss terms. The ChipGAN architecture [35] is an end-to-end GAN model that addresses critical techniques in ink painting, such as blanks, brush strokes, ink tone and spread. Furthermore, an attentional wavelet network [6] utilizes wavelets to capture high-level mood and local details of paintings via the 2-D Haar wavelet transform. Lastly, the Paint-CUT model [36] intelligently creates landscape paintings by utilizing a shuffle attentional residual block and edge enhancement techniques.
@@
For the third type, the output image is guided by textual input information. A novel system [37] transforms classical poetry into corresponding artistic landscape paintings and calligraphic works. Another approach, controlled landscape painting generation (CCLAP) [36], is based on the latent diffusion model (LDM) [38] and comprises two sequential modules: a content generator and a style aggregator. Although this method allows text-based control of the image, it still faces challenges related to the randomness of image generation and limited controllability.
II-B Diffusion-Based Method
@@
DM [39] is a deep learning framework primarily utilized for image processing and computer vision tasks. Fundamentally, DM work by corrupting the training data by continuously adding Gaussian noise, and then learning to recover the data by reversing this noise process. DM can generate detailed images from textual descriptions and are also effective for image restoration, image drawing, text-to-image, and image-to-image tasks [40, 41, 42, 43]. Essentially, by providing a textual description of the desired image, SD can produce a realistic image that aligns with the given description [39]. These models reframe the ”image generation” process into a ”diffusion” process that incrementally removes noise. Starting with random Gaussian noise, the models progressively eliminate it through training until it is entirely removed, yielding an image that closely matches the text description [39, 41]. However, a significant drawback of this approach is the considerable time and memory required, especially for high-resolution image generation [40]. LDM [38] was developed to address these limitations by significantly reducing memory and computational costs. This is achieved by applying the diffusion process within a lower-dimensional latent space instead of the high-dimensional pixel space [40].
@@
The emergence of large text-to-image models that can create visually appealing images from brief descriptive prompts has highlighted the remarkable potential of AI. However, these models encounter challenges such as uneven data availability across specific domains compared to the generalized image-to-text domain. In addition, some tasks require more precise control and guidance than simple prompts can provide. ControlNet [44] addresses these issues by generating high-quality images based on user-provided cues and controls, which can fine-tune performance for specific tasks. Meanwhile, IP-Adapter [45] employs decoupled cross-attention mechanisms for the characteristics of the text and the image. Inspired by these approaches, our study incorporates ControlNet’s additional structural controls and the IP-Adapter’s style guiding the generation of landscape paintings.
III Proposed Method
@@ III-D Training and Inference
$L_{\text{simple}}=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},\bm{c}_{t},\bm{c}_{i},t}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\big{(}\bm{x}_{t},\bm{c}_{t},\bm{c}_{i},t\big{)}\|^{2}.$ (3)
@@
$\begin{split}\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},\bm{c}_{t},\bm{c}_{i},t)=w\bm{\epsilon}_{\theta}(\bm{x}_{t},\bm{c}_{t},\bm{c}_{i},t)+(1-w)\bm{\epsilon}_{\theta}(\bm{x}_{t},t).\end{split}$ (4)
@@ III-E Datasets
Data Collection. Datasets serve as the foundation for the rapid advancement of artificial intelligence, with a few open datasets significantly contributing to this progress. For example, the traditional landscape painting dataset  [31] is the only accessible dataset, but it has limitations in terms of data volume and lacks categorization or identification of the image data. To address this problem, we firstly have curated a collection of digital paintings sourced through websites such as Baidu, artwork websites, photographs from landscape painting books, and digital museum databases. In the following, we enlisted the help of professors and landscape artists to assist in the classification process. All the landscape paintings are classified into three main categories: azure green landscape, ink wash landscape, and light vermilion landscape.
@@
Data Preprocessing. Since the training data for this study should be in $512\times 512$ format, preprocessing the cleaned data is an essential step. First, each image was scaled so that its shorter side was 512 pixels, maintaining the original aspect ratio. For images with an aspect ratio less than 1.5, a $512\times 512$ section was cropped from the center and saved. For images with an aspect ratio greater than 1.5, after cropping the $512\times 512$ section from the center, the center point was moved 256 pixels in both directions along the longer side to crop and save two additional images. If the cropped image was not square, the process was halted, and the current cropped image was discarded. The distribution of the number of pictures is shown in Table I.
TABLE I: Landscape painting type distribution of three distinct styles in our collected dataset. The data comes from Harvard, Smithsonian, Metropolitan, Baidu, and Princeton.
@@
Data Pairing. The data format for this experiment is image-text pairs because it allows the model to learn the association between visual content and descriptive language, enhancing its ability to generate images that align with specific textual descriptions, thereby improving the overall accuracy and relevance of the generated images. Although image-text pairs previously required manual annotation, we now utilize BLIP-2 to automatically generate the corresponding text information for the collected images. Although the text generated by BLIP-2  [46] may be inaccurate, we invited artists who curated descriptive texts specific to landscape paintings. To increase the diversity of text descriptions, we provided multiple expressions for each text and specified the painting type within the descriptions. Finally, we saved the image paths and corresponding text descriptions in JSON files.
III-F Evaluation Metrics
Several metrics are commonly used to assess how closely generated images resemble reference images in evaluating image and structural similarity. This document explains the meanings, formulas and applications of six key metrics: the learned perceptual image patch similarity (LPIPS) [47], gram matrix [48], histogram similarity [49], chamfer match score [50], hausdorff distance [51], and contour match score [52].
@@ III-F Evaluation Metrics
Histogram Similarity (Bhattacharyya Distance). The histogram similarity is a metric for evaluating the similarity of color distributions of two images and is widely used in image processing and computer vision. By comparing the color histograms of the images, the similarity of the images in terms of color can be determined.The bhattacharyya Distance  [53] is a commonly used histogram similarity metric, especially for similarity calculation of probability distributions.
@@
IV Experiment and Analysis
IV-A Generative Controllability
@@
IV-B Qualitative Analysis
To accurately and objectively assess the effectiveness of the proposed LPGen model in generating high-quality landscape paintings, we conduct a comprehensive comparative analysis against several current state-of-the-art methods, including Reference Only [44], Double ControlNet  [44], and Lora  [54].
@@ IV-C Quantitative analysis
TABLE II: Quantitative comparison of the proposed LPGen with several state-of-the-art models. LPIPS evaluates perceptual similarity, gram matrix (GM) measures texture correlations for style similarity, and histogram similarity (HS) assesses color distribution. chamfer match score (CMS i) focuses on average edge similarity, hausdorff distance (HD) emphasizes worst-case edge similarity, and contour match score (CMS ii) evaluates overall shape similarity.
Methods GM $\downarrow$ HS $\downarrow$ LPIPS $\downarrow$ CMS  i $\downarrow$ HD $\downarrow$ CMS ii $\downarrow$
Reference Only [44] 9.21e-06 0.89 0.71 -0.13 212.25 5.76
Double ControlNet [44] 5.45e-06 0.81 0.60 -0.10 209.88 8.38
Lora [54] 3.43e-06 0.75 0.52 -0.10 182.28 7.74
LPGen (Ours) 3.40e-06 0.72 0.55 -0.15 154.54 3.24
@@
The landscape paintings generated by the proposed method were quantitatively compared with those generated by Reference Only, Double ControlNet, and Lora. Each model generated 15 images for each pair of style and structure, resulting in a total of 60 landscape painting images. For each model, the highest values for six metrics of the generated images were recorded: gram matrix similarity, histogram similarity, LPIPS, chamfer match score, hausdorff distance, and contour match score. The scores of the different diffusion models are shown in Table II.
@@
In Table II, the first three metrics, the gram matrix, histogram similarity, and LPIPS, are used to analyze the structural similarity between model outputs and reference images. The table shows that LPGen performs best with respect to the gram matrix, suggesting that LPGen’s generated images have the most similar texture style to the reference image. Furthermore, LPGen performs best with respect to histogram similarity, suggesting that LPGen-generated images have the most similar color distribution to the reference image. For structural similarity, LPGen excels in both gram matrix and histogram similarity, while Lora performs best in LPIPS due to the structure controller. To repetitive fine-tuning, Lora performs best in terms of LPIPS, indicating that Lora’s generated images are most perceptually similar to the reference image. Taking these factors into account, LPGen is the best model for overall structural Tsimilarity.
@@
In Table II, the last three metrics, the chamfer match score, hausdorff distance, and contour match score, are used to analyze the style similarity between different model outputs and the reference image, mainly focusing on edges and contours. Obviously, LPGen performs best with respect to the chamfer match score, suggesting that LPGen’s generated images have the most similar edges to the reference image. Meanwhile, LPGen performs best with respect to the hausdorff distance, suggesting that LPGen’s generated images are closest to the reference image regarding edge similarity. Furthermore, LPGen performs best with respect to the contour match score, indicating that LPGen’s generated images have the most similar contour shapes to the reference image. Considered comprehensively, the landscape paintings generated by the proposed model have the most similar edges and contour shapes to the reference image, making it the best model for style similarity.
IV-D Visual Assessment